# Artificial barriers for stochastic differential equations and for construction of boundary-preserving schemes

### v2

@@ Abstract.
We develop the novel method of artificial barriers for scalar stochastic differential equations (SDEs) and use it to construct boundary-preserving numerical schemes for strong approximations of scalar SDEs, possibly with non-globally Lipschitz drift and diffusion coefficients, whose state-space is either bounded or half-bounded. The idea of artificial barriers is to augment the SDE with artificial barriers outside the state-space to not change the solution process, and then apply a boundary-preserving numerical scheme to the resulting reflected SDE (RSDE). This enables us to construct boundary-preserving numerical schemes with the same strong convergence rate as the strong convergence rate of the numerical scheme for the corresponding RSDE. Based on the method of artificial barriers, we construct two boundary-preserving schemes that we call the Artificial Barriers Euler–Maruyama (ABEM) scheme and the Artificial Barriers Euler–Peano (ABEP) scheme, respectively. We provide numerical experiments for the ABEM scheme and the numerical results agree with the obtained theoretical results.
@@ 1. Introduction
Stochastic differential equations (SDEs) are used to model phenomena arising in, for example, physics, mathematical finance, mathematical biology and epidemic modelling [11, 16, 20, 22, 32]. Some SDEs are confined to a strict subset, referred to as the domain of the SDE, of the target-space. Classical schemes for SDEs are known to leave the domains of SDEs [42]. In this article, we develop the method of artificial barriers and construct, based on this new concept, strongly convergent numerical approximations of SDEs that are confined to the domain of the considered SDE.
@@
with some deterministic initial value $X(0)=x_{0}$, where $B$ is a standard Brownian motion and $T\in(0,\infty)$. We assume that the drift coefficient $f$ and diffusion coefficient $g$ are Lipschitz continuous functions on an open set $D=(a,b)\subset\mathbb{R}$ (see Assumption 1) chosen such that, provided that $x_{0}\in D$, $X(t)\in D$ for every $t\in(0,T]$ almost surely (see Assumption 2). Observe that we assume that $X(t)$ cannot hit the boundary $\partial D=\{a,b\}$ of the domain, which can be characterised by Feller’s classification of unattainable boundaries (see, for example, Chapter $15$ in [16]). To cover both bounded and half-bounded cases, we let $a\in\mathbb{R}$ and $b\in\mathbb{R}\cup\{\infty\}$ be such that $a<b$. We call the case $b=\infty$ the one-finite-boundary case and the case $b\in(a,\infty)$ the two-finite-boundaries case. The case $a=-\infty$ and $b\in(a,\infty)$ is treated in the same way. We do not consider $D=\mathbb{R}$ as the classical Euler–Maruyama (EM) scheme is boundary-preserving in this case.
@@
We say that a numerical scheme $\hat{X}$ for (1) is boundary-preserving if $\hat{X}(t)\in D$ for every $t\in[0,T]$ almost surely. Numerical schemes that may hit the boundary, but not exceed them, are referred to as non-strict boundary-preserving. On one hand, in certain applications, the domain of a model has physical significance. Numerical approximations that leave the physical domain of a model can be interpreted as non-physical. Models defined in a domain show up, for example, in the study of population dynamics [11, 18] and of phase-separation dynamics [2], defined in $[0,1]$, and in the study of temperature dynamics (for example, the classical heat equation), defined in $[0,\infty)$. On the other hand, a numerical scheme that leaves the physical domain of a model will naturally be affected by forces that do not influence the model. Let us consider an example in the setting of this paper: If the solution of the SDE (1) only takes values in $D=(0,1)$, then modifying $f$ and $g$ outside $D=(0,1)$ will not impact the solution. Modifying $f$ and $g$ outside $D=(0,1)$ can, however, influence a numerical approximation that can leave the domain $D=(0,1)$. We can avoid such discrepancy between the exact solution and numerical approximations by constructing and studying boundary-preserving numerical schemes.
@@
The motivation for the current work is that most previously proposed and studied boundary-preserving schemes for SDEs have either exploited the Lamperti transform or time splitting procedures (see, for example, [30]) to confine the numerical approximations to the domain of the SDE. We also mention [4, 5, 13, 14, 19, 37, 43] for some other proposed approaches to construct boundary-preserving schemes. The Lamperti transform is used to transform the considered SDE into another SDE with additive noise (see, for example, [29]), that is easier to solve or numerically integrate. Then, under some assumptions, optimal convergence rates for numerical schemes for SDEs with additive noise can be transferred to the original SDE [7, 42]. Traditionally, numerical schemes utilising the Lamperti transform have combined it with some classical scheme (for example, the EM scheme or the backward EM scheme) for SDEs (see, for example, [1, 31, 44]). Recently, authors (see [17, 42]) have combined the Lamperti transform with time splitting methods to construct efficient numerical approximations to certain types of SDEs. Lamperti-based schemes do, however, require high regularity assumptions (for example, differentiability) on the drift coefficient $f$ and on the diffusion coefficient $g$ and require the existence of an inverse of the Lamperti transform to be well-defined. For instance, $L^{p}(\Omega)$-convergence rate of $1$, for all $p\geq 1$, was obtained in [42] for a two-finite-boundaries setting with drift coefficient $f\in C^{2}$ and diffusion coefficient $g\in C^{3}$, and $g$ satisfying non-integrability conditions near the boundary points. See, for example, Theorem 22 for the definition of $L^{p}(\Omega)$-convergence used in this work. In addition, the Lamperti transform is intrinsically problem-specific which makes the implementation of Lamperti-based schemes problem-specific.
@@
The objective of this paper is to construct boundary-preserving schemes for the SDE (1) that do not depend on the Lamperti transform and for which optimal $L^{p}(\Omega)$-convergence can be obtained with lower regularity assumptions on the drift and diffusion coefficients $f$ and $g$. To this end, we develop the novel method of artificial barriers and combine it with modifications of two classical schemes for reflected stochastic differential equations (RSDEs) to construct non-Lamperti-based boundary-preserving schemes for the SDE (1). We refer readers to [8, 34, 36, 38, 41] for works on the theory of RSDEs and to [9, 24, 25, 26, 33, 39, 40] for numerical approximations of RSDEs. Our proposed Artificial Barriers Euler–Maruyama (ABEM) scheme also has the advantages that it is less involved to implement (no need to compute auxiliary functions such as the Lamperti transform and its inverse), is it less sensitive to initial values close to the boundary and, to the best of our knowledge, it is the first implementable boundary-preserving numerical scheme that can approximate solutions of SDEs in the one-finite-boundary and in the two-finite-boundary case in one unified framework.
@@
  • We prove that the proposed method of artificial barriers exhibits optimal $L^{p}(\Omega)$-convergence order, for every $p\geq 2$, when combined with an optimal order $L^{p}(\Omega)$-convergent scheme for reflected SDEs. Moreover, we also obtain almost sure pathwise convergence as a corollary (see Corollary 23). In particular, we prove that artificial barriers combined with a modified projected EM scheme (referred to as the ABEM scheme) exhibits $L^{p}(\Omega)$-convergence order of $1/2-$ for every $p\geq 2$ (see Theorem 25) and combined with a modified Euler–Peano scheme (referred to as the ABEP scheme) exhibits $L^{p}(\Omega)$-convergence order of $1/2$ for every $p\geq 2$ (see Theorem 28).
@@ 2. Setting
[…]
@@
Note that $D$ depends on the choices of $f$ and $g$. Let $a\in\mathbb{R}$ and $b\in\mathbb{R}\cup\{\infty\}$, with $a<b$, be the boundary points of $D$: $\partial D=\{a,b\}$. The case $a=-\infty$ and $b\in\mathbb{R}$ is treated in the same way. We say that a stochastic process $\left(X(t)\right)_{t\in[0,T]}$ is a (strong) solution of (2) if the corresponding Itô integral equation
@@
is satisfied, almost surely, for every $t\in[0,T]$, where the second integral is an Itô integral. We refer to [32] for a well-know textbook on SDEs.
@@
The drift coefficient $f:D\to\mathbb{R}$ and diffusion coefficient $g:D\to\mathbb{R}$ are Lipschitz continuous functions and we denote by $L_{f}=\operatorname{Lip}(f)$ and $L_{g}=\operatorname{Lip}(g)$ the Lipschitz constants of $f$ and $g$ on $D$.
@@
Assumption 1 only assumes that the drift and diffusion coefficient functions are Lipschitz continuous functions on the domain $D$. If $|D|<\infty$, then $f$ and $g$ (view as functions $\mathbb{R}\to\mathbb{R}$) can be chosen to be locally Lipschitz functions with superlinear growth. See Section 7 for a possible choice of superlinearly growing $f$ and $g$.
@@
(4) $|g(x)|\leq C(L_{g})\left(1+|x|\right),\ \forall x\in D.$
@@
The drift coefficient $f:D\to\mathbb{R}$ and diffusion coefficient $g:D\to\mathbb{R}$ are chosen such that the boundary points $a$ and $b$ are unattainable by solutions $X$ of (2); that is, we assume that ${\mathbb{P}}(X(t)\in D,\ \forall t\in[0,T])=1$.
@@
for every $p\geq 2$, and that only takes values in $D$
@@ 3. Skorokhod maps and Reflected SDEs
Here we introduce Skorokhod maps and define reflected SDEs in terms of Skorokhod maps. Shortly put, Skorokhod maps (in our case) are constructed to map any path in $\mathcal{D}[0,T]$ to a path in $\mathcal{D}[0,T]$ confined to a subset of $\mathbb{R}$. Reflected SDEs, also known as SDEs with boundary conditions, are essentially SDEs that are only allowed to take values in a certain subset of $\mathbb{R}$, if the process hits the boundary of the subset then the process is reflected.
@@ 3.1. Skorokhod maps
First, we introduce and provide some properties of Skorokhod maps that underlies the theory of reflected SDEs discussed in the next section. In essence, a Skorokhod map takes, in our case, a function in $\mathcal{D}[0,T]$ and outputs another function in $\mathcal{D}[0,T]$ that is confined to the domain under consideration. We say that a Skorokhod map $\Gamma$ is associated with a domain $A$ if $\Gamma(\psi)(t)\in A$ for all $\psi\in\mathcal{D}[0,T]$ and for all $t\in[0,T]$. We refer to Skorokhod maps associated with half-bounded domains $[\alpha,\infty)$, for $\alpha\in\mathbb{R}$, as one-sided Skorokhod maps and we refer to Skorokhod maps associated with bounded domains $[\alpha,\beta]$, for $\alpha<\beta$ and $|\alpha|+|\beta|<\infty$, as two-sided Skorokhod maps. One-sided Skorokhod maps were first considered in [38]. Moreover, an explicit formula for one-sided Skorokhod maps (see Definition 3 below) has been known since long (see, for example, [40]). An explicit formula for two-sided Skorokhod maps (see Definition 5 below) was only later derived (see [23]). We are interested in Skorokhod maps for the purpose of constructing implementable numerical schemes and for proving $L^{p}(\Omega)$-convergence. Note, however, that explicit formulas for Skorokhod maps do not guarantee that numerical schemes based on Skorokhod maps can be implemented (see, for example, the ABEP scheme in Section 6). Certain schemes based on Skorokhod maps can, however, be implemented. Most importantly, some piecewise constant schemes combined with a Skorokhod map can be implemented (see, for example, the ABEM scheme in Section 5).
@@
The one-sided Skorokhod map $\Gamma^{0}:\mathcal{D}[0,T]\to\mathcal{D}[0,T]$ associated with the domain $[0,\infty)$ is the map defined by
(5) $\Gamma^{0}(\psi)(t)=\psi(t)+\sup_{s\in[0,t]}\max\{0,-\psi(s)\}$
for $\psi\in\mathcal{D}[0,T]$ and for $t\in[0,T]$. The one-sided Skorokhod map $\Gamma^{\alpha}:\mathcal{D}[0,T]\to\mathcal{D}[0,T]$, $\alpha\in\mathbb{R}$, associated with the domain $[\alpha,\infty)$ is defined by
(6) $\Gamma^{\alpha}(\psi)(t)=\Gamma^{0}(\psi)(t)+\alpha$
for $\psi\in\mathcal{D}[0,T]$ and for $t\in[0,T]$.
@@
for any $\psi_{1},\ \psi_{2}\in\mathcal{D}[0,T]$ and for any $r\in[0,T]$, and the following growth bound in the supremum norm
(8) $\sup_{t\in[0,r]}|\Gamma^{\alpha}(\psi)(t)|\leq 2\sup_{t\in[0,r]}|\psi(t)|+|\alpha|,$
for any $\psi\in\mathcal{D}[0,T]$ and for any $r\in[0,T]$.
@@
The two-sided Skorokhod map $\Gamma^{0,\alpha}:\mathcal{D}[0,T]\to\mathcal{D}[0,T]$, for $\alpha>0$, associated with the domain $[0,\alpha]$ is the map defined by
@@
for $\psi\in\mathcal{D}[0,T]$, where $\Gamma^{0}$ is the one-sided Skorokhod map associated with the domain $[0,\infty)$ defined in equation (5) and where the map $\Lambda^{\alpha}:\mathcal{D}[0,T]\to\mathcal{D}[0,T]$ is defined by
$\Lambda^{\alpha}(\psi)(t)=\psi(t)-\sup_{s\in[0,t]}\left(\min\left(\max(0,\psi(s)-\alpha),\inf_{u\in[s,t]}\psi(u)\right)\right).$
@@
Similarly to $\Gamma^{0}$, $\Gamma^{0,\alpha}$ satisfies a Lipschitz estimate in the supremum norm and a bound in the supremum norm that we state in the following lemma.
@@
for any $\psi_{1},\ \psi_{2}\in\mathcal{D}[0,T]$ and for any $r\in[0,T]$, and is bounded in the supremum norm
@@
We are interested in the Skorokhod map $\Gamma$ associated with $\operatorname{cl}D=(a,b)\cup\{a,b\}$ and the family of Skorokhod maps $\{\Gamma_{N}\}_{N\in\mathbb{N}}$ each associated with $\operatorname{cl}D_{N}=\partial D_{N}\cup D_{N}$. The maps $\Gamma$ and $\Gamma_{N}$ can be expressed in terms of $\Gamma^{0}$ and $\Gamma^{0,\alpha}$ as
(12) $\Gamma(\psi)(t)=\begin{cases}\Gamma^{a}(\psi)(t),\text{ if }b=\infty,\\ \Gamma^{0,(b-a)}(\psi)(t)+a,\text{ if }|b|<\infty,\end{cases}$
@@
(13) $\Gamma_{N}(\psi)(t)=\begin{cases}\Gamma^{a+1/N}(\psi)(t),\text{ if }b=\infty,\\ \Gamma^{0,(b-a-2/N)}(\psi)(t)+(a+1/N),\text{ if }|b|<\infty,\end{cases}$
@@
$\sup_{t\in[0,r]}|\Gamma(\psi)(t)|\leq\begin{cases}2\sup_{t\in[0,r]}|\psi(t)|+|a|,\text{ if }b=\infty,\\ \max(|a|,|b|),\text{ if }|b|<\infty,\end{cases}$
@@
(15) $\sup_{N\in\mathbb{N}}\sup_{t\in[0,r]}|\Gamma_{N}(\psi)(t)|\leq\begin{cases}2\sup_{t\in[0,r]}|\psi(t)|+|a|+1,\text{ if }b=\infty,\\ \max(|a|,|b|),\text{ if }|b|<\infty,\end{cases}$
@@
$\sup_{t\in[0,r]}|\Gamma_{N}(\psi)(t)|=\sup_{t\in[0,r]}|\Gamma^{a+1/N}(\psi)(t)|\leq 2\sup_{t\in[0,r]}|\psi(t)|+\left|a+\frac{1}{N}\right|\leq 2\sup_{t\in[0,r]}|\psi(t)|+|a|+1.$

### v3

@@ Abstract.
We develop the novel method of artificial barriers for scalar stochastic differential equations (SDEs) and use it to construct boundary-preserving numerical schemes for strong approximation of scalar SDEs, possibly with non-globally Lipschitz drift and diffusion coefficients, whose state-space is either bounded or half-bounded. The idea of artificial barriers is to augment the SDE with artificial barriers outside the state-space to not change the solution process, and then apply a boundary-preserving numerical scheme to the resulting reflected SDE (RSDE). This enables us to construct boundary-preserving numerical schemes that achieve the same strong convergence rate as the corresponding RSDE scheme. Based on the method of artificial barriers, we construct two boundary-preserving schemes that we call the Artificial Barriers Euler–Maruyama (ABEM) scheme and the Artificial Barriers Euler–Peano (ABEP) scheme, respectively. We provide numerical experiments for the ABEM scheme and the numerical results agree with the obtained theoretical results.
@@ 1. Introduction
Stochastic differential equations (SDEs) are used to model phenomena arising in, for example, physics, mathematical finance, mathematical biology and epidemic modelling [15, 20, 24, 26, 40]. Some SDEs are confined to a strict subset, referred to as the domain of the SDE, of the target-space. Classical schemes for SDEs are known to leave the domains of SDEs [50]. In this article, we develop the method of artificial barriers and construct, based on this new concept, strongly convergent numerical approximations of solutions of SDEs that are confined to the domain of the considered SDE.
@@
with some deterministic initial value $X(0)=x_{0}$, where $B$ is a standard Brownian motion and $T\in(0,\infty)$. We assume that the drift coefficient $f$ and diffusion coefficient $g$ are Lipschitz continuous functions on an open set $D=(a,b)\subset\mathbb{R}$ (see Assumption 1) such that, provided that $x_{0}\in D$, $X(t)\in D$ for every $t\in(0,T]$, almost surely (see Assumption 2). Observe that we assume that $X(t)$ cannot hit the boundary $\partial D=\{a,b\}$ of the domain, which can be characterised by Feller’s classification of unattainable boundaries (see, for example, Chapter $15$ in [20]). To cover both bounded and half-bounded cases, we let $a\in\mathbb{R}$ and $b\in\mathbb{R}\cup\{\infty\}$ be such that $a<b$. We call the case $b=\infty$ the one-finite-boundary case and the case $b\in(a,\infty)$ the two-finite-boundaries case. The case $a=$-$\infty$ and $b\in(a,\infty)$ is treated in the same way. We do not consider $D=\mathbb{R}$ as the classical Euler–Maruyama (EM) scheme is boundary-preserving in this case.
@@
We say that a numerical scheme $\hat{X}$ for (1) is boundary-preserving if $\hat{X}(t)\in D$ for every $t\in[0,T]$, almost surely. Numerical schemes that may hit the boundary, but not exceed them, are referred to as non-strict boundary-preserving. On one hand, in certain applications, the domain of a model has physical significance. Numerical approximations that leave the physical domain of a model can be interpreted as non-physical. Models defined in a domain show up, for example, in the study of population dynamics [15, 22] and of phase-separation dynamics [2], defined in $[0,1]$, and in the study of temperature dynamics (for example, the classical heat equation), defined in $0,\infty)$. On the other hand, a numerical scheme that leaves the physical domain of a model will naturally be affected by forces that do not influence the model. Let us consider an example in the setting of this paper: If the solution of the SDE (1) only takes values in $D=(0,1)$, then modifying $f$ and $g$ outside $D=(0,1)$ will not impact the solution. Modifying $f$ and $g$ outside $D=(0,1)$ can, however, influence a numerical approximation that can leave the domain $D=(0,1)$. We can avoid such discrepancy between the exact solution and numerical approximations by constructing and studying boundary-preserving numerical schemes.
Let us discuss previous approaches to construct boundary-preserving numerical schemes. Early boundary-preserving schemes were typically implicit schemes. See, for example, [1, 11, 12, 39, 45] for some works based on implicit techniques. Implicit schemes, however, tend to be computationally slower than explicit ones. The main focus in this work is explicit schemes, but implicit schemes could also be used within the developed framework. Another common approach to construct boundary-preserving schemes is to transform the solution to be approximated to another process that is easier to construct such schemes for. Typical examples of such transformations include the Lamperti transform (see, for example, [37]) and the logarithmic transform [52]. Time splitting methods [38], truncation techniques [28, 35], and approximation by geometric Brownian motion [4, 5] have also successfully been employed to construct boundary-preserving numerical schemes. An issue with the mentioned approaches, except schemes based on truncation, is that they requires regularity assumptions on the coefficient functions. Truncation schemes are the most related schemes to what we proposed in this work. We also mention [6, 7, 17, 18, 23, 51] for some other proposed approaches to construct boundary-preserving schemes.
@@
Let us discuss the Lamperti transform a bit more in detail. Under some assumptions, optimal convergence rates for numerical schemes for SDEs with additive noise can be transferred to these SDEs using the Lamperti transform [9, 50]. Recently, authors (see [30, 31]) have constructed higher-order boundary-preserving numerical schemes using the Lamperti transform. Traditionally, numerical schemes utilising the Lamperti transform have combined it with some classical scheme (for example, the EM scheme or the backward EM scheme) for SDEs (see, for example, [1, 39, 52]). Recently, authors (see [21, 50]) have combined the Lamperti transform with time splitting methods to construct efficient numerical approximations the solutions to certain types of SDEs. Lamperti-based schemes do, however, require high regularity assumptions (for example, differentiability) on the drift coefficient $f$ and on the diffusion coefficient $g$ and require the existence of an inverse of the Lamperti transform to be well-defined. For instance, $L^{p}(\Omega)$-convergence rate of $1$, for all $p\geq 1$, was obtained in [50] for an SDE with bounded domain with drift coefficient $f\in C^{2}$ and diffusion coefficient $g\in C^{3}$, and $f$ and $g$ satisfying some technical conditions near the boundary points. See, for example, Theorem 22 for the definition of $L^{p}(\Omega)$-convergence used in this work. In addition, the Lamperti transform is intrinsically problem-specific which makes the implementation of Lamperti-based schemes problem-specific.
@@
The objective of this paper is to construct boundary-preserving schemes for the SDE (1) that do not depend on the Lamperti transform and for which optimal $L^{p}(\Omega)$-convergence can be obtained with lower regularity assumptions on the drift and diffusion coefficients $f$ and $g$. To this end, we develop modified versions of two classical schemes for reflected stochastic differential equations (RSDEs) to construct non-Lamperti-based boundary-preserving schemes for the SDE (1). We refer readers to [10, 42, 44, 46, 49] for works on the theory of RSDEs and to [13, 29, 32, 33, 41, 47, 48] for numerical approximations of RSDEs.
@@
  • We prove that the proposed method of artificial barriers exhibits optimal $L^{p}(\Omega)$-convergence order, for every $p\geq 2$, when combined with an optimal order $L^{p}(\Omega)$-convergent scheme for reflected SDEs. Moreover, we also obtain almost sure pathwise convergence as a corollary (see Corollary 23). In particular, we prove that, for Lipschitz continuous coefficient functions, artificial barriers combined with a modified projected EM scheme (referred to as the ABEM scheme) exhibits $L^{p}(\Omega)$-convergence order of $1/2$- for every $p\geq 2$ (see Theorem 25) and combined with a modified Euler–Peano scheme (referred to as the ABEP scheme) exhibits $L^{p}(\Omega)$-convergence order of $1/2$ for every $p\geq 2$ (see Theorem 28).
  • We numerically verify boundary-preservation and $L^{2}(\Omega)$-convergence of order $1/2$- for the ABEM scheme.
@@ 2. Setting
(2) $\left\{\begin{aligned} &\mathop{}\!\mathrm{d}X(t)=f(X(t))\mathop{}\!\mathrm{d}t+g(X(t))\mathop{}\!\mathrm{d}B(t),\ t\in(0,T],\\ &X(0)=x_{0}\in D,\end{aligned}\right$.
@@
Note that $D$ depends on the choices of $f$ and $g$. Let $a\in\mathbb{R}$ and $b\in\mathbb{R}\cup\{\infty\}$ be such that $a<0<b$, and be the boundary points of $D$: $\partial D=\{a,b\}$. The other cases are treated similarly. We say that a stochastic process $\left(X(t)\right)_{t\in[0,T]}$ is a (strong) solution of (2) if the corresponding Itô integral equation
@@
is satisfied, for every $t\in[0,T]$, almost surely, where the second integral is an Itô integral. We refer to [40] for a well-know textbook on Itô SDEs.
@@
The drift coefficient $f:D\to\mathbb{R}$ and diffusion coefficient $g:D\to\mathbb{R}$ are Lipschitz continuous functions with Lipschitz constants $\operatorname{Lip}(f)$ and $\operatorname{Lip}(g)$, respectively; that is, there exists $\operatorname{Lip}(f),\operatorname{Lip}(f)>0$ such that
$|f(x)-f(y)|\leq\operatorname{Lip}(f)|x-y|,\ \forall x,y\in D,$
and
$|g(x)-g(y)|\leq\operatorname{Lip}(g)|x-y|,\ \forall x,y\in D.$
@@
Assumption 1 only assumes that the drift and diffusion coefficient functions are Lipschitz continuous functions on the domain $D$. If $|D|<\infty$, then $f$ and $g$ (view as functions $\mathbb{R}\to\mathbb{R}$) can be locally Lipschitz functions with superlinear growth. See Section 7 for a possible choice of superlinearly growing $f$ and $g$.
@@
(4) $|g(x)|\leq|g(0)|+\operatorname{Lip}(g)|x|,\ \forall x\in D.$
We introduce
$L_{f}=\max\left(|f(0)|,\operatorname{Lip}(f)\right),$
to have one common bound for $\operatorname{Lip}(f)$ and for the linear growth constant in (3). For the same reason, we let
$L_{g}=\max\left(|g(0)|,\operatorname{Lip}(g)\right).$
@@
The drift coefficient $f:D\to\mathbb{R}$ and diffusion coefficient $g:D\to\mathbb{R}$ are such that the boundary points $a$ and $b$ are unattainable by solutions $X$ of (2); that is, we assume that ${\mathbb{P}}(X(t)\in D,\ \forall t\in[0,T])=1$.
@@
for every $p\geq 2$, and satisfying
@@ 3. Skorokhod maps and Reflected SDEs
Here we introduce Skorokhod maps and define reflected SDEs (RSDEs) in terms of Skorokhod maps. Shortly put, Skorokhod maps (in our case) are constructed to map any path in $\mathcal{D}[0,T]$ to a path in $\mathcal{D}[0,T]$ confined to a subset of $\mathbb{R}$. RSDEs, also known as SDEs with boundary conditions, are essentially SDEs that are only allowed to take values in a certain subset of $\mathbb{R}$, if the process hits the boundary of the subset then the process is reflected.
@@ 3.1. Skorokhod maps
First, we introduce and provide some properties of Skorokhod maps underlying the theory of RSDEs discussed in the next section. In essence, a Skorokhod map takes, in our case, a function in $\mathcal{D}[0,T]$ and outputs another function in $\mathcal{D}[0,T]$ that is confined to the domain under consideration. We say that a Skorokhod map $\Gamma$ is associated with a domain $A$ if $\Gamma(\psi)(t)\in A$ for all $\psi\in\mathcal{D}[0,T]$ and for all $t\in[0,T]$. We refer to Skorokhod maps associated with half-bounded domains $\alpha,\infty)$, for $\alpha\in\mathbb{R}$, as one-sided Skorokhod maps and we refer to Skorokhod maps associated with bounded domains $\alpha,\beta]$, for $\alpha<\beta$ and $|\alpha|+|\beta|<\infty$, as two-sided Skorokhod maps. One-sided Skorokhod maps were first considered in [46]. Moreover, an explicit formula for one-sided Skorokhod maps (see Definition 3 below) has been known since long (see, for example, [48]). An explicit formula for two-sided Skorokhod maps (see Definition 5 below) was only later derived (see [27]). We are interested in Skorokhod maps for the purpose of constructing implementable numerical schemes and for proving $L^{p}(\Omega)$-convergence. Note, however, that explicit formulas for Skorokhod maps do not guarantee that numerical schemes based on Skorokhod maps can be implemented (see, for example, the ABEP scheme in Section 6). Certain schemes based on Skorokhod maps can, however, be implemented. Most importantly, some piecewise constant schemes combined with a Skorokhod map can be implemented (see, for example, the ABEM scheme in Section 5).
@@
The one-sided Skorokhod map $\Gamma^{0}:\mathcal{D}[0,T]\to\mathcal{D}[0,T]$ associated with the domain $0,\infty)$ is the map defined by
(5) $\Gamma^{0}(\psi)(t)=\psi(t)+\sup_{s\in[0,t]}\max\{0,-\psi(s)\},\ \psi\in\mathcal{D}[0,T],\ t\in[0,T].$
The one-sided Skorokhod map $\Gamma^{\alpha}:\mathcal{D}[0,T]\to\mathcal{D}[0,T]$, $\alpha\neq 0$, associated with the domain $\alpha,\infty)$ is defined by
(6) $\Gamma^{\alpha}(\psi)(t)=\Gamma^{0}(\psi-\alpha)(t)+\alpha,\ \psi\in\mathcal{D}[0,T],\ t\in[0,T].$
@@
for every $\psi_{1},\ \psi_{2}\in\mathcal{D}[0,T]$ and for every $r\in[0,T]$, and the following growth bound in the supremum norm
(8) $\sup_{t\in[0,r]}|\Gamma^{\alpha}(\psi)(t)|\leq 2\sup_{t\in[0,r]}|\psi(t)|+3|\alpha|,$
for every $\psi\in\mathcal{D}[0,T]$ and for every $r\in[0,T]$.
@@
The two-sided Skorokhod map $\Gamma^{0,\alpha}:\mathcal{D}[0,T]\to\mathcal{D}[0,T]$, for $\alpha>0$, associated with the domain $0,\alpha]$ is the map defined by
@@
where $\Gamma^{0}$ is the one-sided Skorokhod map associated with the domain $0,\infty)$ defined in equation (5) and where the map $\Lambda^{\alpha}:\mathcal{D}[0,T]\to\mathcal{D}[0,T]$ is defined by
$\Lambda^{\alpha}(\psi)(t)=\psi(t)-\sup_{s\in[0,t]}\left(\min\left(\max(0,\psi(s)-\alpha),\inf_{u\in[s,t]}\psi(u)\right)\right),\ \psi\in\mathcal{D}[0,T].$
@@
Similarly to $\Gamma^{\alpha}$, $\Gamma^{0,\alpha}$ satisfies a Lipschitz estimate in the supremum norm and a bound in the supremum norm that we state in the following lemma.
@@
for every $\psi_{1},\ \psi_{2}\in\mathcal{D}[0,T]$ and for every $r\in[0,T]$, and is bounded in the supremum norm
@@
We are interested in the Skorokhod map $\Gamma$ associated with $\operatorname{cl}D=(a,b)\cup\{a,b\}$ and the family of Skorokhod maps $\{\Gamma_{N}\}_{N\in\mathbb{N}}$ each associated with $\operatorname{cl}D_{N}=\partial D_{N}\cup D_{N}$. The maps $\Gamma$ and $\Gamma_{N}$ can be expressed in terms of $\Gamma^{\alpha}$ and $\Gamma^{0,\alpha}$ as
(12) $\Gamma(\psi)(t)=\begin{cases}\Gamma^{a}(\psi)(t),\text{ if }b=\infty,\\ \Gamma^{0,(b-a)}(\psi-a)(t)+a,\text{ if }|b|<\infty,\end{cases}$
@@
(13) $\Gamma_{N}(\psi)(t)=\begin{cases}\Gamma^{a+1/N}(\psi)(t),\text{ if }b=\infty,\\ \Gamma^{0,(b-a-2/N)}\left(\psi-(a+1/N)\right)(t)+(a+1/N),\text{ if }|b|<\infty,\end{cases}$
@@
$\sup_{t\in[0,r]}|\Gamma(\psi)(t)|\leq\begin{cases}2\sup_{t\in[0,r]}|\psi(t)|+3|a|,\text{ if }b=\infty,\\ \max(|a|,|b|),\text{ if }|b|<\infty,\end{cases}$
@@
(15) $\sup_{N\in\mathbb{N}}\sup_{t\in[0,r]}|\Gamma_{N}(\psi)(t)|\leq\begin{cases}2\sup_{t\in[0,r]}|\psi(t)|+3|a|+3,\text{ if }b=\infty,\\ \max(|a|,|b|),\text{ if }|b|<\infty,\end{cases}$
@@
$\sup_{t\in[0,r]}|\Gamma_{N}(\psi)(t)|=\sup_{t\in[0,r]}|\Gamma^{a+1/N}(\psi)(t)|\leq 2\sup_{t\in[0,r]}|\psi(t)|+3\left|a+\frac{1}{N}\right|\leq 2\sup_{t\in[0,r]}|\psi(t)|+3|a|+3.$